OpenStudy (anonymous):

A simple pendulum on the surface of Earth is found to undergo 11 complete small-amplitude oscillations in 7.8s. Find the pendulum's length.

OpenStudy (anonymous):

The restoring force of a pendulum \[-mg \sin \theta = ma\] \[a=-g \sin \theta \] For small angles, \[\sin \theta = \theta\] so \[a=\frac{\mathrm{d}^2v}{\mathrm{d}t^2}=\ell \frac{\mathrm{d}^2\theta}{\mathrm{d}t^2}=-g\theta\] giving the simple diff eq \[\frac{\mathrm{d}^2\theta}{\mathrm{d}t^2}+\frac{g}{\ell}\theta=0\] with the solution \[\theta=\theta_{max} \sin(\omega t)\] \[ \text{angular frequency} \ = \omega = \sqrt{\frac{g}{\ell}}\] \[ \text{Period}=T=\frac{2\pi}{\omega} = 2\pi \sqrt{\frac{\ell}{g}}\] \[t=7.8s \\ n = \text{number of oscillations}=11\] \[T = \frac{\text{time}}{\text{number of oscillations}}=\frac{t}{n}\] \[\text{equate the two expressions for period, and solve for} \ \ \ell\] ^_^